### v1

# Active Learning with Variational Quantum Circuits for Quantum Process Tomography

Abstract
Quantum process tomography (QPT), used for reconstruction of an unknown quantum process from measurement data, is a fundamental tool for the diagnostic and full characterization of quantum systems. It relies on querying a set of quantum states as input to the quantum process. Previous works commonly use a straightforward strategy to select a set of quantum states randomly, overlooking differences in informativeness among quantum states. Since querying the quantum system requires multiple experiments that can be prohibitively costly, it is always the case that there are not enough quantum states for high-quality reconstruction. In this paper, we propose a general framework for active learning (AL) to adaptively select a set of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used variational quantum circuits (VQCs) to perform the QPT task and integrate our AL algorithms into the query step. We design and evaluate three various types of AL algorithms: committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost. Additionally, we provide a guideline for selecting algorithms suitable for different scenarios. Numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline method that selects a set of quantum states randomly. Moreover, these results suggest that active learning based approaches are applicable to other complicated learning tasks in large-scale quantum information processing.
Corresponding author.
Introduction
With recent advances in quantum computing and machine learning, their intersection has emerged as an increasingly attractive area of research (Biamonte et al. 2017; Schuld, Sinayskiy, and Petruccione 2015). On one hand, quantum computing enhances the efficiency of machine learning algorithms by exploiting quantum advantage (Dunjko, Taylor, and Briegel 2016; Wiebe, Kapoor, and Svore 2014). On the other hand, machine learning methodologies and techniques are employed to help solving problems in quantum computation and quantum information (Carleo and Troyer 2017; Mitarai et al. 2018). In this paper, we apply active learning (AL) to solve a challenging problem of quantum process tomography (QPT).
QPT, used for reconstruction of an unknown quantum process from measurement data, is a fundamental tool for the diagnostic and full characterization of quantum systems (Chuang and Nielsen 1997). Standard quantum process tomography (SQPT) relies on an informationally complete set of quantum states as input, which inevitably leads to an algorithmic complexity in terms of the number of measurements and classical post-processing that scales exponentially with the system size (D’Ariano and Presti 2001). Due to these limitations, SQPT has only been experimentally implemented up to three qubits (Krinner et al. 2020; Muñoz-Gil et al. 2020). However, with the rapid advancement of quantum computing hardware, there is a growing demand for scalable QPT schemes (Guo et al. 2019; Wu et al. 2021).
Inspired by machine learning techniques, various methods have been proposed to alleviate the exponential scaling problem. Some statistical methods such as maximum-likelihood estimation (MLE) (Smolin, Gambetta, and Smith 2012), Bayesian estimation (Granade, Combes, and Cory 2016), and projected-least-squares (PLS) (Surawy-Stepney et al. 2022) techniques have been adapted for QPT. A major challenge in current applications is the experimental and computational cost of estimating large-scale systems. A class of methods partially address this challenge, including compressed-sensing QPT (CS-QPT) (Rodionov et al. 2014; Flammia et al. 2012), tensor network QPT (TN-QPT) (Torlai et al. 2023), but at the cost of assuming specific structures of the unknown quantum process. Recently, quantum machine learning combined with variational quantum circuits (VQCs) has received extensive attention (Abbas et al. 2021; Torlai et al. 2018). Another promising class of methods focuses on reconstructing a unitary quantum process by inverting the dynamics using variational algorithms (Xue et al. 2022; Carolan et al. 2020).
Indeed, although these methods employ different techniques for reconstruction, they all involve a crucial step: query, which requires selecting a subset of quantum states from the informationally complete set as input of the unknown process. It is worth noting that quantum states with greater informativeness typically improving reconstruction more efficiently. However, previous works commonly use a straightforward strategy to select the subset randomly, overlooking differences in informativeness among quantum states. Since querying the quantum system requires multiple experiments that can be prohibitively costly, it is always the case that there are not enough quantum states for high-quality reconstruction, especially in large-scale quantum systems. Building on this observation, this work does not aim to replace these methods but to propose a general framework that further enhances them by optimizing the query step. Specifically, we focus on the informativeness of quantum states, aiming to select a subset of quantum states that improves the reconstruction most efficiently.
AL is a promising machine learning paradigm that helps reduce the annotation costs by intelligently selecting a subset of informative samples for labeling (Settles 2009). With the aim of reducing experiment costs, AL has been extended from computer science to physics and applied successfully in various subfields such as condensed matter physics (Zhang et al. 2019; Yao et al. 2020), high-energy physics (Caron et al. 2019), and quantum information (Ding et al. 2020; Lange et al. 2023; Dutt et al. 2023). Inspired by this, we first utilize the differences in infromativeness among quantum states and propose a general framework for AL to adaptively select a subset of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. Additionally, we design and evaluate three various types of AL algorithms with the aim of offering choices suited to different scenarios, including committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost (Kumar and Gupta 2020).
Contributions.
The main contributions of this work are summarized as follows. First, we propose a general framework for AL that enhances previous methods by optimizing the crucial query step. To the best of our knowledge, this is the first framework that utilizes the differences in informativeness among quantum states. In particular, we introduce a learning framework that leverages the widely-used VQC to perform the QPT task and integrate our AL algorithms into the query step. Second, we design and evaluate three various types of AL algorithms: committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost. Third, numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline that selects a subset of quantum states randomly. Additionally, these results further indicate the importance of our algorithms, especially in large-scale quantum systems. Lastly, we provide a guideline for selecting algorithms suitable for different scenarios.
Preliminaries
Figure 1: Hybrid quantum-classical procedure of executing and optimizing a variational quantum circuit (VQC).
In this section, we briefly introduce the basic concepts of quantum computation and quantum information that are necessary in this work. Interested readers are recommended to the celebrated textbook by Nielsen and Chuang (Nielsen and Chuang 2001).
Qubit, Quantum State.
Qubit is the fundamental concept of quantum computation and quantum information. Unlike a classical bit is strictly binary, a qubit can be in a superposition of both binary states. A single-qubit quantum state is a linear combination of two basis states, represented as $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ in Dirac notation, where $\alpha,\beta\in\mathbb{C}$ and $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$. Here, $|0\rangle$ and $|1\rangle$ denote the basis states $\left[1,0\right]^{T}$ and $\left[0,1\right]^{T}\in\mathbb{C}^{2}$, respectively. Similarly, an $n$-qubit quantum state can be represented as $|\psi\rangle=\sum_{i=0}^{2^{n}-1}\alpha_{i}|i\rangle$, where $\alpha_{i}\in\mathbb{C}$, $|i\rangle\in\mathbb{C}^{2^{n}}$ for each $i$, and $\sum_{i}\left|\alpha_{i}\right|^{2}=1$.
Unitary Transformation, Pauli Matrices.
An $n$-qubit unitary transformation is defined as an operator $U\in\mathbb{C}^{2^{n}\times 2^{n}}$ that satisfies $U^{\dagger}U=I$, where $U^{\dagger}$ denotes the conjugate transpose of $U$. The operator $U$ transforms the quantum state of the system from $|\psi\rangle$ to $U|\psi\rangle$. Pauli matrices are key examples of such unitary transformations:
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad Y=\begin{pmatrix}0&-\textnormal{i}\\ \textnormal{i}&0\end{pmatrix},\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
Quantum Gate, Quantum Circuit.
Some important quantum gates featured extensively in this work include the Pauli rotations $\left\{R_{P}(\theta)=\textnormal{e}^{-\textnormal{i}\frac{\theta}{2}P}\mid P\in\left\{X,Y,Z\right\}\right\}$, which are exponential forms of the Pauli matrices. Additionally, the controlled-X (CNOT) gate, represented as $\text{CNOT}=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X$, is used to generate quantum entanglement among qubits. A sequence of quantum gates applied to a system of qubits forms a quantum circuit, at the end of which the qubits are measured to obtain classical values.
Figure 2: Framework of active learning with variational quantum circuits for quantum process tomography.
Variational Quantum Circuit.
VQC is a hybrid quantum-classical approach that leverages the strengths of both quantum and classical computation. It is a class of quantum circuits with tunable parameters, optimized for a specific purpose on a classical computer, while the variational circuit itself is executed on a quantum computer, as illustrated in Fig. 1. This approach has the advantage of keeping the quantum circuit depth shallow and hence mitigating noise. Consequently, VQCs have been developed for a wide range of applications, covering essentially all of the applications that researchers had envisaged for quantum computers, and they appear to be the best hope for obtaining quantum advantage (Cerezo et al. 2021).
Approximating Unitary Quantum Process with Variational Quantum Circuits
In this section, we discuss how the design leverages widely-used VQCs to perform the QPT task and provide a detailed explanation of the learning framework.
Model
Unknown Quantum Process.
Let $\mathcal{U}$ be an unknown unitary quantum process. It is reasonable to compile it into a variational quantum circuit (Khatri et al. 2019; Xue et al. 2022). More specifically, we can approximate $\mathcal{U}$ using a VQC denoted by $\mathcal{C}(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the parameters to be optimized. Consequently, all the information about $\mathcal{U}$ is encoded in the parameters $\boldsymbol{\theta}$, enabling us to systematically reconstruct $\mathcal{U}$ from these parameters using a classical computer.
Variational Quantum Circuit Ansatz.
The ansatz of VQC is shown in Fig. 2. It is composed of interlaced layers using single-qubit Pauli rotation gates and two-qubit CNOT gates, designed to quickly generate entanglement between qubits. Thus, it makes it possible to approximate complicated quantum processes. Each two-qubit layer is counted as a depth and each single-qubit layer contains three rotational gates ($R_{z}$, $R_{y}$ and $R_{z}$) on each qubit. For such a circuit with $n$ qubits and $d$ depths, the total number of parameters is $3n(d+1)$.
Learning Framework
The learning framework encompasses three key components: quantum state preparation, query, and training, as shown in Fig. 2.
Quantum State Preparation.
A quantum state is prepared by applying a unitary transformation $U$ to the initial state: $|0\rangle^{\otimes n}\rightarrow U|0\rangle^{\otimes n}$. Specifically, the unitary transformation $U$ is selected from tensor products of single-qubit Pauli operators $\left\{I,X,Y,Z\right\}^{\otimes n}$. We generate an informationally complete set of quantum states, denoted as $D=\left\{\left|\psi_{1}\right\rangle,\left|\psi_{2}\right\rangle,\ldots,\left|\psi_{N}\right\rangle\right\}$, where $N=4^{n}$.
Query.
We refer to the unknown unitary process $\mathcal{U}$ as an oracle that can be queried. When a quantum state $|\psi\rangle$ is queried to the oracle, it returns the state $|\psi^{\mathrm{ideal}}\rangle=\mathcal{U}|\psi\rangle$ as its true label. This procedure allows us to construct a training dataset.
Training.
Given a training dataset $D_{l}=\left\{\left(\left|\psi_{1}\right\rangle,\left|\psi_{1}^{\mathrm{ideal}}\right\rangle\right),\left(\left|\psi_{2}\right\rangle,\left|\psi_{2}^{\mathrm{ideal}}\right\rangle\right),\ldots,\left(\left|\psi_{m}\right\rangle,\left|\psi_{m}^{\mathrm{ideal}}\right\rangle\right)\right\}$, the loss function is defined as the mean square error:
$\mathcal{L}(\boldsymbol{\theta})$ $=\frac{1}{m}\sum_{i=1}^{m}\left\|\left|\psi_{i}(\boldsymbol{\theta})\right\rangle-\left|\psi_{i}^{\mathrm{ideal}}\right\rangle\right\|^{2}$
$=\frac{1}{m}\sum_{i=1}^{m}\bigg{(}\left\langle\psi_{i}(\boldsymbol{\theta})\mid\psi_{i}(\boldsymbol{\theta})\right\rangle+\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}^{\mathrm{ideal}}\right\rangle$
$\quad-\left\langle\psi_{i}(\boldsymbol{\theta})\mid\psi_{i}^{\mathrm{ideal}}\right\rangle-\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\boldsymbol{\theta})\right\rangle\bigg{)}$
$=\frac{2}{m}\sum_{i=1}^{m}\left[1-\mathrm{Re}\left(\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\boldsymbol{\theta})\right\rangle\right)\right],$ (1)
where $\left|\psi_{i}\left(\boldsymbol{\theta}\right)\right\rangle=\mathcal{C}\left(\boldsymbol{\theta}\right)\left|\psi_{i}\right\rangle$ denotes the predicted label. The inner product on the right-hand side can be efficiently computed using a generalized SWAP-test algorithm on a quantum computer (Xue et al. 2022).
To minimize the loss function, we employ a gradient-based optimization method to iteratively update the parameters $\boldsymbol{\theta}$. Using the chain rule, the gradient is computed as
$\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\theta_{i}}=\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}},$ (2)
where $\mathcal{O}$ represents the observable corresponding to the generalized SWAP test. The first term, $\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}$, can be computed classically once the value of $\mathcal{O}(\boldsymbol{\theta})$ is obtained from the quantum computer. The second term, $\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}$, is computed on the quantum computer using the parameter-shift rule (Mitarai et al. 2018):
$\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}=\frac{1}{2}\left[\mathcal{O}(\boldsymbol{\theta}_{i}^{+})-\mathcal{O}(\boldsymbol{\theta}_{i}^{-})\right],$ (3)
where $\boldsymbol{\theta}_{i}^{\pm}$ denotes the parameter vector $\boldsymbol{\theta}$ with the $i$th parameter shifted by ±$\pm\frac{\pi}{2}$.
Method
In this section, we first analyze the inefficiencies in the crucial query step. Then, we propose a general framework for AL that utilizes the differences in informativeness among quantum states to optimize this step. Finally, we design various AL algorithms with the aim of offering choices suited to different scenarios.
Active Learning
Considering that querying a quantum system is prohibitively costly since it requires multiple experiments, our goal is to learn the parameters $\boldsymbol{\theta}$ with error $\epsilon$ from training dataset $\{(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\}_{i=1}^{m}$, while minimizing $m$. However, in the query step, it is common to select a subset of quantum states randomly. This straightforward strategy typically requires a large number of labeled quantum states to train a VQC to achieve satisfactory performance. In fact, each quantum state contains different informativeness, with those having greater informativeness typically improving training more efficiently. This random selection strategy does not take this into account, resulting in training inefficiencies. To overcome this limitation, we focus on the informativeness of quantum states, aiming to select a subset of quantum states that improves the training most efficiently.
Acquisition Function.
AL is a machine learning paradigm that promises to reduce annotation costs by intelligently selecting a subset of informative samples for labeling (Settles 2009). The selection of such a subset relies on a so-called acquisition function, which is key to AL. Specifically, the acquisition function evaluates the informativeness of each unlabeled sample and ranks them. The highest-ranking sample is then selected for labeling and subsequently added to the training dataset for updating the model. This process is repeated, gradually increasing the number of labeled samples.
Active Learning Framework.
Our framework for AL is illustrated in Fig. 2 and can be briefly described as follows:
  • Initialize: Initialize an unlabeled pool $D_{u}$ with the informationally complete set of quantum states, an empty labeled pool $D_{l}$, and a base model VQC.
  • Evaluate: Use the acquisition function to evaluate and rank each quantum state in the unlabeled pool $D_{u}$.
  • Select: Select the highest-ranking quantum state $|\psi^{\star}\rangle$.
  • Query: Query the unknown unitary process $\mathcal{U}$ to obtain the label for the selected quantum state.
  • Add: Add the newly labeled quantum state to the labeled pool $D_{l}$.
  • Update: Update the model with the newly incorporated labeled pool $D_{l}$.
The process follows an iterative procedure in which quantum states are labeled sequentially to improve the model performance. At each AL step, the most informative quantum state is selected based on the acquisition function. This cycle is repeated until the labeling budget is exhausted or a satisfactory model performance is achieved.
Based on various acquisition functions, current AL methods can be broadly classified into committee-based, uncertainty-based and diversity-based (Kumar and Gupta 2020). With the aim of offering choices suited to different scenarios, we design various AL algorithms to each category as follows.
Query-by-Committee
Query-by-committee (QBC) is a widely used AL approach (Freund et al. 1997; Burbidge, Rowland, and King 2007). Its basic idea is to build a committee of models from existing labeled data, and then selects the sample for which the committee disagree most to label.
We construct a committee of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, which are initialized with different parameters and trained on $D_{l}$. In each AL step, each member of the committee $\boldsymbol{\Theta}$ casts its vote based on their learned quantum process representations $\{\mathcal{C}(\boldsymbol{\theta}_{1}),\mathcal{C}(\boldsymbol{\theta}_{2}),\ldots,\mathcal{C}(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\}$. The committee select the quantum state for which all models disagree most on the predictions $\{|\psi(\boldsymbol{\theta}_{1})\rangle,|\psi(\boldsymbol{\theta}_{2})\rangle\},\ldots,|\psi(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\rangle\}$. A common strategy to calculate the level of disagreement is the variance. More precisely, for each unlabeled quantum state $|\psi\rangle\in D_{u}$, we compute the variance of the $n_{\mathrm{VQC}}$ individual predictions and then select the quantum state with the maximum variance, that is,
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}d\left(|\psi(\boldsymbol{\theta}_{k})\rangle,|\psi(\boldsymbol{\Theta})\rangle\right),$ (4)
where $|\psi({\boldsymbol{\Theta}})\rangle\propto\sum_{k=1}^{n_{\mathrm{VQC}}}|\psi({\boldsymbol{\theta}_{k}})\rangle$, and $d$ represents the Euclidean distance. The pseudo-code for QBC is shown in Algorithm 1.
Algorithm 1 Query-by-Committee (QBC)
Input: Labeled quantum states set $D_{l}$, unlabeled quantum states set $D_{u}$, number of committee members $n_{\mathrm{VQC}}$
Output: Selected quantum state $|\psi^{\star}\rangle$ with the maximum variance
1:  Construct a committee of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
2:  for $|\psi\rangle\in D_{u}$ do
3:     for $k=1$ to $n_{\mathrm{VQC}}$ do
4:        $|\psi(\boldsymbol{\theta}_{k})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta}_{k})|\psi\rangle$
5:     end for
6:     Calculate the variance of $\{|\psi(\boldsymbol{\theta}_{k})\rangle\}_{k=1}^{n_{\mathrm{VQC}}}$ using Eq. (4)
7:  end for
Expected Model Change Maximization
Expected model change maximization (EMCM) is also a popular AL approach that aims to select the sample resulting in the largest change to the current model (Cai, Zhang, and Zhou 2013). The model change is measured as the difference between the current model parameters and the updated parameters after training with the enlarged training dataset. Inspired by Stochastic Gradient Descent (SGD) update rule, the change is estimated as the gradient of the loss with respect to an unlabeled sample.
Specifically, we employ SGD to update the parameters $\boldsymbol{\theta}$. The parameters $\boldsymbol{\theta}$ are updated iteratively according to the negative gradient descent of the loss $\ell(\boldsymbol{\theta})$ with respect to each training data $(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\in D_{l}$:
$\boldsymbol{\theta}:=\boldsymbol{\theta}-\alpha\nabla_{\boldsymbol{\theta}}\ell_{|\psi_{i}\rangle}(\boldsymbol{\theta})\quad\text{for }i=1,2,\ldots,m,$ (5)
where $\alpha$ denotes the learning rate.
Here, we consider the update rule in AL process. Assume an unlabeled quantum state $|\psi\rangle\in D_{u}$, which is then labeled as $|\psi^{\mathrm{ideal}}\rangle$ and added to training dataset. The loss function on the enlarged training dataset $D_{l}^{+}=D_{l}\cup(|\psi\rangle,|\psi^{\mathrm{ideal}}\rangle)$ then becomes:
$\mathcal{L}_{D_{l}^{+}}=\frac{1}{|D_{l}^{+}|}\left[\sum_{|\psi_{i}\rangle\in D_{l}}\left\||\psi_{i}\rangle-|\psi_{i}^{\mathrm{ideal}}\rangle\right\|^{2}+\underbrace{\left\||\psi\rangle-|\psi^{\mathrm{ideal}}\rangle\right\|^{2}}_{=:\ell_{|\psi\rangle}(\boldsymbol{\theta})}\right].$ (6)
Thus, the model obtained by minimizing the loss function is changed due to the inclusion of the new training data $(|\psi\rangle,|\psi^{\mathrm{ideal}}\rangle)$. We measure the model change as the parameter change using the gradient of the loss at $|\psi\rangle$. The derivative of the loss $\ell_{|\psi\rangle}(\boldsymbol{\theta})$ with respect to the parameters $\theta_{i}$ at $|\psi\rangle$ is formulated as:
$\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\theta_{i}}$ $=\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}$
$=\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\left[\frac{1}{2}\left(\mathcal{O}(\boldsymbol{\theta}_{i}^{+})-\mathcal{O}(\boldsymbol{\theta}_{i}^{-})\right)\right],$ (7)
where $\mathcal{O}(\boldsymbol{\theta})=\mathrm{Re}\left(\left\langle\psi^{\mathrm{ideal}}\mid\psi(\boldsymbol{\theta})\right\rangle\right)$. Since the true label $|\psi^{\mathrm{ideal}}\rangle$ is actually unknown before querying in practice, we construct an ensemble $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$ to estimate the prediction $\{|\psi(\boldsymbol{\theta}_{1})\rangle,|\psi(\boldsymbol{\theta}_{2})\rangle,\ldots,|\psi(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\rangle\}$, and use the expected model change to approximate the true model change, that is:
$\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\theta_{i}}$ $=\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}\frac{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}{\partial\theta_{i}}$
$=\frac{1}{2n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}\left[\hat{\mathcal{O}}_{k}(\boldsymbol{\theta}_{i}^{+})-\hat{\mathcal{O}}_{k}(\boldsymbol{\theta}_{i}^{-})\right],$ (8)
where $\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})=\mathrm{Re}\left(\left\langle\psi(\boldsymbol{\theta}_{k})\mid\psi(\boldsymbol{\theta})\right\rangle\right)$ represents an estimate of $\mathcal{O}$.
The goal of the sampling strategy is to select the quantum state that could maximally change the current model. Therefore, the acquisition function can be formulated as:
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\left\|\nabla_{\boldsymbol{\theta}}\ell_{|\psi\rangle}(\boldsymbol{\theta})\right\|.$ (9)
The pseudo-code for EMCM is shown in Algorithm 2.
Algorithm 2 Expected Model Change Maximization (EMCM)
Input: Labeled quantum states set $D_{l}$, unlabeled quantum states set $D_{u}$, base model $\mathcal{C}(\boldsymbol{\theta})$ trained on $D_{l}$, number of ensemble models $n_{\mathrm{VQC}}$
Output: Selected quantum state $|\psi^{\star}\rangle$ with the maximum gradient norm
1:  Construct an ensemble of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
2:  for $|\psi\rangle\in D_{u}$ do
3:     $|\psi(\boldsymbol{\theta})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta})|\psi\rangle$
4:     for $k=1$ to $n_{\mathrm{VQC}}$ do
5:        $|\psi(\boldsymbol{\theta}_{k})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta}_{k})|\psi\rangle$
6:     end for
7:     Calculate the gradient norm $\left\|\nabla_{\boldsymbol{\theta}}\ell_{|\psi\rangle}(\boldsymbol{\theta})\right\|$ using Eq.(8)
8:  end for
Greedy Sampling
Instead of finding the most informative sample based on the learned model, as in QBC and EMCM, greedy sampling (GS) is a model-free approach that selects the sample based on its geometric characteristics in the feature space (Yu and Kim 2010). Its basic idea is to greedily select an unlabeled sample that is located far from the previously labeled samples, thereby ensuring that the selected samples are scattered across the entire input space rather than concentrating in a small local region. Clearly, this approach allows for selecting a set of samples at the beginning of the AL process, making it easy to implement by avoiding the need to evaluate each unlabeled sample and update the model at each AL step.
For each unlabeled quantum state $|\psi\rangle\in D_{u}$, we compute its distance to all labeled quantum states in $D_{l}$. The quantum state that maximizes the its distance to the labeled quantum state set is selected. Formally,
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\underset{|\psi_{i}\rangle\in D_{l}}{\operatorname{min}}d(|\psi\rangle,|\psi_{i}\rangle),$ (10)
where $d$ represents the Euclidean distance. The pseudo-code for GS is shown in Algorithm 3.
Algorithm 3 Greedy Sampling (GS)
Input: Labeled quantum states set $D_{l}$, unlabeled quantum states set $D_{u}$
Output: Selected the quantum state $|\psi^{\star}\rangle$ with the maximum of the minimum distance
1:  for $|\psi\rangle\in D_{u}$ do
2:     for $|\psi_{i}\rangle\in D_{l}$ do
3:        Calculate the distance $d(|\psi\rangle,|\psi_{i}\rangle)$
4:     end for
5:     Find the minimum distance $\underset{|\psi_{i}\rangle\in D_{l}}{\operatorname{min}}d(|\psi\rangle,|\psi_{i}\rangle)$
6:  end for
Numerical Experiments
In this section, we validate our algorithms through numerical simulations on the quantum processes of randomly generated circuits. Specifically, we conduct simulations with up to five qubits to evaluate the scalability of our algorithms.
Experimental Settings
Unknown Unitary Process.
We consider the quantum process of randomly generated circuits, as shown in Fig. 3. The circuit begins and ends with Hadamard gates applied to each qubit. Each two-qubit layer is counted as a depth, including controlled-phase (CZ) gates alternating between odd and even depths, and randomly chosen single-qubit gate ($T$, $\sqrt{X}$, or $\sqrt{Y}$). It is noted that such randomly generated quantum circuits are hard for efficient simulation on a classical computer (Boixo et al. 2018; Bouland et al. 2019).
Figure 3: Randomly generated quantum circuit
Baseline and Performance Metrics.
We consider a baseline random (RAND) that selects a subset of quantum states randomly. Additionally, we use similarity between the target process $\mathcal{U}$ and the approximate reconstruction $\mathcal{C}(\boldsymbol{\theta})$ to measure performance, defined as
$\textnormal{similarity}\left(\mathcal{U},\mathcal{C}(\boldsymbol{\theta})\right)=1-\frac{\|\mathcal{C}(\boldsymbol{\theta})-\mathcal{U}\|_{F}}{2\|\mathcal{U}\|_{F}}.$
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The $\text{similarity}(\mathcal{U},\mathcal{C}(\boldsymbol{\theta}))=1$ means $\|\mathcal{C}(\theta)-\mathcal{U}\|_{F}=0$, in which case $\mathcal{U}$ can be perfectly reconstructed from $\mathcal{C}(\boldsymbol{\theta})$.
To better evaluate our algorithms, we define an additional metric called improvement, which is the ratio of the similarity achieved by our algorithms to that achieved by RAND, with the same number of labeled quantum states.
Other Experimental Parameters.
We conduct simulations with different numbers of qubits to evaluate the scalability of our algorithms. For an $n$-qubit quantum process ($n$ ranges from 2 to 5), we set the VQC depth to 3, 5, 7 and 8 to learn the corresponding quantum processes. The number of committee members in QBC and the number of ensemble models in EMCM is empirically set to be 6. To avoid random fluctuation, each experiment is repeated 100 times and the averaged metrics are reported.
Experimental Results
Figure 4: The comparison results with the metric of average similarity.
Figure 5: The comparison results with the metric of average improvement.
Similarity Metric.
The results related to similarity are shown in Fig. 4. We observe the following.
  • Generally, as the number of labeled quantum states increased, all four algorithms achieved better performance. This is intuitive, as more training samples typically result in a more reliable model.
  • QBC, EMCM, and GS converges faster than RAND, each achieving higher similarity with significantly fewer labeled quantum states. This clearly demonstrates that our algorithms are highly effective in selecting informative quantum states, thereby improving the reconstruction efficiency.
  • QBC and EMCM significantly outperformed RAND across all numbers of labeled quantum states. GS initially performs worse than RAND when only a few quantum states are labeled. However, as the number of labeled quantum states increases, GS eventually outperforms RAND. A possible explanation for the phenomena is as follows. At the initial steps, GS might select some outliers to maximize the distance from the previously labeled quantum states, which leads to a decrease in performance. Since the number of outliers is usually very small, their impact diminishes as more quantum states are labeled. It is worth noting that as the system size increases, reconstructing the quantum process requires a significantly larger number of labeled quantum states, which helps mitigate the impact of outliers. Therefore, GS remains a promising approach, especially in large-scale quantum systems.
Improvement Metric.
The results related to improvement are shown in Fig. 5. We observe the following.
  • When $n$ ranges from 2 to 4, the improvement of QBC and EMCM initially increases and then decreases. This is because, with sufficient labeled quantum states, even randomly selected quantum states can lead to well-trained models. Meanwhile, the improvement of GS initially decreases, which is consistent with our previous analysis, as GS may select outliers in the initial steps.
  • When $n$ increases to 5, the improvement of all our algorithms does not exhibit a significant decreasing trend. This is because, as the system size increases, RAND requires more labeled quantum states to train a high quality model. Despite a large number of quantum states already being labeled, which supports good performance for our algorithms, this is still insufficient for RAND, resulting in consistently poor performance. Thus, our algorithms are crucial in large-scale quantum systems, where RAND always fails to obtain enough labeled quantum states to train a reliable model.
  • When $n$ increases further, the maximum achievable improvement of all our algorithms increases, which further demonstrates the effectiveness of our algorithms in large-scale systems.
Guideline for Algorithms Selection.
It is evident that QBC and EMCM achieved better performance than GS. However, both QBC and EMCM are model-based algorithms that require evaluating unlabeled quantum states and updating the model at each AL step. Additionally, they require training multiple models, which is computationally expensive. In contrast, GS is easy to implement and computationally inexpensive, as it simply selects the quantum state based on its geometric characteristics. Considering performance, query cost, and computational cost, we can conclude that:
  • In small-scale quantum systems or scenarios where query cost severely limits the number of queries, QBC and EMCM are generally preferable as they offer reliable performance.
  • In large-scale quantum systems, if computational cost is acceptable compared to query cost, QBC and EMCM are more suitable choices. Conversely, if computational cost is a major concern, GS is a more suitable choice.
Conclusion
In this work, we apply AL to solve a challenging problem of QPT. We propose a general framework for AL to adaptively select a subset of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. We design and evaluate three various types of AL algorithms and provide a guideline for selecting algorithms suitable for different scenarios. Numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline that selects a subset of quantum states randomly. Moreover, these results indicate the importance of our algorithms in large-scale quantum information processing scenarios.
For future work, it will be of great interest to extend AL to other costly tasks in quantum information processing, such as shadow tomography, quantum detection tomography (QDT), quantum sensing, and quantum control (Gebhart et al. 2023).
Acknowledgements
[…]